OpenStudy (batabooche):

State the horizontal asymptote of the rational function. f(x) = quantity five x plus one divided by quantity nine x minus two.

OpenStudy (welshfella):

5x + 1 ------ 9x - 2

OpenStudy (welshfella):

what does the value of the function approach when x approaches positive or negative infinity?

OpenStudy (batabooche):

No..I'm not that good at these

OpenStudy (welshfella):

or if the word infinity confuses you a bit very large positive or negative numbers.

OpenStudy (welshfella):

Ok so x = 1000 if we plug this into the function we get 5001 ---- 8998 what simple fraction is that close to?

OpenStudy (welshfella):

we can take the top to be 5000 and bottom to be 9000 That simplifies to what?

OpenStudy (welshfella):

do you see how i got 5001? = 5x + 1 x = 1000 so we have 5(1000) + 1 = 5001.

OpenStudy (welshfella):

so what is 5000/9000 as a simple fraction - divide both top and bottom by 1000 . What do you get?

OpenStudy (welshfella):

5000 / 1000 = ?

OpenStudy (welshfella):

5 isn't it?

OpenStudy (welshfella):

I can't help you if you don't participate in the post. I could give you a direct answer but that would be against the Code of Conduct. And you wouldn't learn anything that way.